OpenStudy (anonymous):

kay..herz the problem: x+4y-z=6 3x+2y+3z=16 2x-y+z=3

OpenStudy (anonymous):

whoops..tats confusing...wait...

OpenStudy (saifoo.khan):

Who wants to get pissed?

OpenStudy (anonymous):

kay..herz the problem: x+4y-z=6 3x+2y+3z=16 2x-y+z=3

OpenStudy (mimi_x3):

Oh gawd, im already pissed off now :@ Saifi, im not in a good mood today, please.

OpenStudy (anonymous):

dunt worry..if he does, ima report him!

OpenStudy (mimi_x3):

Which method first ? Cramer or Substitution ?

OpenStudy (anonymous):

what ever i easier!

OpenStudy (saifoo.khan):

Third one.

OpenStudy (saifoo.khan):

But why are u pissed? did your iphone broke? Did u get an A in Maths or what?

OpenStudy (anonymous):

plz..if ur not gonna help seafood..GO AWAY!!!

OpenStudy (mimi_x3):

I'm depressed T_T

OpenStudy (anonymous):

kay go on H.Mini

OpenStudy (mimi_x3):